# NovelHopQA: Diagnosing Multi-Hop Reasoning Failures in Long Narrative Contexts

### v1

@@ Abstract
Current large language models (LLMs) struggle to answer questions that span tens of thousands of tokens, especially when multi-hop reasoning is involved. While prior benchmarks explore long-context comprehension or multi-hop reasoning in isolation, none jointly vary context length and reasoning depth in natural narrative settings. We introduce NovelHopQA, the first benchmark to evaluate k1–4 hop QA over 64k–128k-token excerpts from 83 full-length public-domain novels. A keyword-guided pipeline builds hop-separated chains grounded in coherent storylines. We evaluate six state-of-the-art (SOTA) models and apply oracle-context filtering to ensure all questions are genuinely answerable. Human annotators validate both alignment and hop depth. We noticed consistent accuracy drops with increased hops and context length, even in frontier models—revealing that sheer scale does not guarantee robust reasoning. Our failure mode analysis highlights common breakdowns, such as missed final-hop integration and long-range drift. NovelHopQA offers a controlled diagnostic setting to stress-test multi-hop reasoning at scale.
@@
Abhay Guptathanks: Lead Author  Michael Luthanks: Lead Investigator Kevin Zhuthanks: Senior AuthorSean O’Brien33footnotemark: 3Vasu Sharma33footnotemark: 3 Algoverse AI Research abhaygupta1266@gmail.com, kevin@algoverse.us
@@ 1 Introduction
    Empirical hop-depth study: evaluations on six SOTA models trace accuracy decay along both axes.
@@ 2 Related Work
Architectural, retrieval, and memory methods for long contexts. To process longer inputs, sparse-attention and recurrence-based architectures—Longformer, BigBird, Transformer-XL, and LongRoPE—scale attention and positional encodings to tens or hundreds of thousands of tokens Beltagy et al. (2020); Zaheer et al. (2021); Dai et al. (2019); Ding et al. (2024). Retrieval-augmented generation and external-memory approaches boost performance when evidence is scattered Lewis et al. (2021); Wu et al. (2022). Stress-test challenges like “Lost in the Middle” and NeedleBench highlight positional and retrieval brittleness in passages Liu et al. (2023b); Li et al. (2024b), while BABILong probes reasoning limits with synthetic million-token haystacks Kuratov et al. (2024). Although these advances surface key failure modes, they do not explore how reasoning depth interacts with very long contexts in natural prose.
@@ 3.1 Source Corpus
Hop o1 4o 4o-mini Gemini 2.5 P Gemini 2.0 F Gemini 2.0 FL Avg. 1 95.90 95.60 92.30 96.80 93.10 90.90 94.10 2 95.50 95.40 91.80 96.50 92.80 90.30 93.72 3 95.20 95.10 91.30 96.30 92.40 90.00 93.38 4 94.80 94.90 90.90 96.20 92.10 89.60 93.08 Avg. 95.35 95.25 91.58 96.45 92.60 90.20 93.57
Table 1: Accuracy (%) of each model on NovelHopQA when evaluated using the original golden context.
Figure 1: Keyword-guided paragraph-chaining pipeline used to build NovelHopQA. See Appendix F for a full example showing multi-hop evolution across four refinement stages.
@@ 3.2 Salient Keyword Filtering
For each of the 83 novels, we prompt GPT-4o-mini OpenAI (2024a) to suggest five “anchor” keywords—characters, locations, or objects central to the plot (see Appendix H for prompt). If any keyword appears fewer than 50 times in the text, we discard and re-sample that anchor, repeating up to seven times to ensure five high-frequency anchors.
Context Hop o1 4o 4o-mini Gemini 2.5 P Gemini 2.0 F Gemini 2.0 FL Avg. 64k 1 92.51 90.12 75.49 92.34 87.37 82.53 86.73 2 87.66↓4.85 84.25↓5.87 74.77↓0.72 87.84↓4.50 77.02↓10.35 71.39↓11.14 80.48↓6.25 3 84.99↓2.67 81.34↓2.91 73.14↓1.63 85.12↓2.72 74.25↓2.77 70.05↓1.34 78.13↓2.35 4 82.15↓2.84 78.47↓2.87 68.04↓5.10 82.45↓2.67 71.76↓2.49 65.33↓4.72 74.69↓3.44 96k 1 90.35 88.83 72.25 90.12 82.26 78.44 83.71 2 85.88↓4.47 82.67↓6.16 67.44↓4.81 86.03↓4.09 74.02↓8.24 67.04↓11.40 77.18↓6.53 3 83.41↓2.47 80.41↓2.42 66.97↓0.47 83.71↓2.32 73.38↓0.64 66.05↓0.99 75.66↓1.52 4 80.68↓2.73 76.92↓3.91 65.59↓1.38 80.98↓2.73 70.26↓3.12 62.81↓3.24 72.87↓2.79 128k 1 88.76 86.95 70.03 89.10 81.77 75.31 81.99 2 84.33↓4.43 80.52↓6.43 63.95↓6.08 84.70↓4.40 69.13↓12.64 62.21↓13.10 74.14↓7.85 3 81.92↓2.41 78.03↓2.92 62.95↓1.00 82.20↓2.50 68.78↓1.35 62.07↓0.14 72.66↓1.48 4 78.80↓3.12 74.64↓3.31 61.18↓1.77 78.55↓3.65 67.32↓1.46 57.39↓4.68 69.65↓3.01
@@ 3.4 Multi-Hop Context Chaining & Incremental QA Generation
For each book and hop depth $H\in\{1,2,3,4\}$, we assemble contexts and QA pairs as follows (see Appendix H for all prompts):
@@
    Paragraph exclusivity: Remove each selected paragraph from the pool to prevent reuse. If no matching paragraph is found after seven attempts, abort the chain and restart with a fresh anchor.
@@ 3.5 Golden-Context Filtering
To verify answerability, we evaluate all six models on the original golden contexts used to generate each QA pair. As shown in Table 1, all models score above 90% on average, confirming the validity of most questions. We discard any question missed by any model in the final dataset used in Section 5. Removal counts are reported in Appendix B.
@@ 3.6 Irrelevant and No-Context Sanity Check
To validate that the questions require contextual reasoning to be solved, we evaluated 800 QA pairs—100 per hop—under irrelevant and no context settings. Removing the context yields low accuracies, suggesting that the tested models are typically unable to answer correctly without contextual grounding. This helps ensure the dataset reflects reasoning, not recall. Full results are in Appendix E.
@@ 4 Human Evaluation
Ten undergraduate validators each annotated 260 examples—40 from the 1- and 2-hop sets, and 90 from the 3- and 4-hop sets. They rated Alignment, measuring how well each QA pair matched its source context, and judged Hop Match, assessing whether the answer required exactly $H$ reasoning steps. See Appendix C for detailed results and Appendix G for the evaluation form.
@@ 5 Results and Discussion
We evaluate six models on NovelHopQA using chain-of-thought prompts: o1 OpenAI (2024c), Gemini 2.5 Pro DeepMind (2025b), GPT-4o OpenAI (2024a), GPT-4o-mini OpenAI (2024a), Gemini 2.0 Flash, and Gemini 2.0 Flash Lite DeepMind (2025a). Table 3 summarizes model accuracy across three context lengths (64k, 96k, 128k) and four hop depths (1–4).
Impact of hop depth. All models exhibit consistent performance degradation as hop depth increases. On average, accuracy drops roughly 12 points from 1-hop to 4-hop at 64k context length. Even reasoning models like Gemini 2.5 Pro and o1 see steady declines with more complex multi-step questions, highlighting the challenge of compositional reasoning at scale.
Impact of context length. Longer context lengths also lead to reduced accuracy, though the effect is milder than that of hop count. Across models, 1-hop performance drops about 5 points when moving from 64k to 128k contexts. This suggests that deeper reasoning contributes more to failure than sheer length.
Model comparisons. Reasoning models—Gemini 2.5 Pro and o1—consistently outperform others, often topping each row in Table 3. Gemini 2.5 Pro achieves the highest average accuracy overall, followed closely by o1 and GPT-4o. Mid-sized models like GPT-4o-mini and Gemini Flash Lite perform noticeably worse, especially under 4-hop and 128k settings, where their scores fall into the 60s.
Robustness at scale. Despite large context windows, no model maintains strong performance on the hardest tasks (4-hop at 128k), where even top models dip below 80%. These results affirm that long-context capacity alone is not enough—robust multi-hop reasoning remains an open challenge. A deeper analysis as to why models fail is provided in Appendix D.
@@ 6 Conclusion
NovelHopQA is the first benchmark to vary both context length (64k–128k) and hop depth $H\in\{1,2,3,4\}$ in long-context QA. Human validation confirms quality, and models show accuracy drops along both axes. These results highlight that larger context windows aren’t enough—multi-hop reasoning remains a core challenge. Code and data will be released upon publication.
@@ 8 Ethics Statement
We describe our dataset construction process in Section 3, and include all prompt templates in Appendix H. All model generations were obtained using publicly available APIs. Specifically, we used the Azure OpenAI API for GPT-4o, GPT-4o-mini OpenAI (2024a), and o1 OpenAI (2024c), and the Google Vertex API for Gemini 2.0 Flash, Flash Lite DeepMind (2025a), and Gemini 2.5 Pro DeepMind (2025b). All models were queried using CoT prompts, and their outputs were graded using GPT-4.1 OpenAI (2024b) with CoT-based evaluation prompts. We plan to release the dataset, prompts, and model outputs upon publication to support replication and further research.
@@ Appendix E Irrelevant and No Context Evaluation
Interpretation. This experiment validates the integrity of NovelHopQA by confirming that models are not simply memorizing QA pairs seen during pretraining. Accuracy remains near-zero when relevant context is removed, demonstrating that our questions are novel and context-dependent. These findings strengthen confidence that model performance on NovelHopQA reflects actual reading comprehension and not artifact exploitation or memorization.
Appendix F 4-Hop QA Evolution Example
@@
Appendix G Human Evaluation Form Example
@@
Appendix H Prompt Templates

### v2

@@ Abstract
Current large language models (LLMs) struggle to answer questions that span tens of thousands of tokens, especially when multi-hop reasoning is involved. While prior benchmarks explore long-context comprehension or multi-hop reasoning in isolation, none jointly vary context length and reasoning depth in natural narrative settings. We introduce NovelHopQA, the first benchmark to evaluate k1–4 hop QA over 64k–128k-token excerpts from 83 full-length public-domain novels. A keyword-guided pipeline builds hop-separated chains grounded in coherent storylines. We evaluate seven state-of-the-art (SOTA) models and apply oracle-context filtering to ensure all questions are genuinely answerable. Human annotators validate both alignment and hop depth.We additionally present retrieval-augmented generation (RAG) evaluations to test model performance when only selected passages are provided instead of the full context. We noticed consistent accuracy drops with increased hops and context length, even in frontier models—revealing that sheer scale does not guarantee robust reasoning. Our failure mode analysis highlights common breakdowns, such as missed final-hop integration and long-range drift. NovelHopQA offers a controlled diagnostic setting to test multi-hop reasoning at scale. All code and datasets are available at: https://novelhopqa.github.io.
@@
Abhay Gupta1thanks: Lead Author  Michael Lu2 Kevin Zhu1Sean O’Brien1Vasu Sharma3 1Algoverse AI Research  2University of California, Berkeley  3Meta FAIR Lab abhay@algoverse.us, kevin@algoverse.us
@@ 1 Introduction
    Empirical hop-depth study: evaluations on seven SOTA models trace accuracy decay along both axes.
@@ 2 Related Work
Architectural, retrieval, and memory methods for long contexts. To process longer inputs, sparse-attention and recurrence-based architectures—Longformer, BigBird, Transformer-XL, and LongRoPE—scale attention and positional encodings to tens or hundreds of thousands of tokens Beltagy et al. (2020); Zaheer et al. (2021); Dai et al. (2019); Ding et al. (2024). RAG and external-memory approaches boost performance when evidence is scattered Lewis et al. (2021); Wu et al. (2022). Stress-test challenges like “Lost in the Middle” and NeedleBench highlight positional and retrieval brittleness in passages Liu et al. (2023b); Li et al. (2024b), while BABILong probes reasoning limits with synthetic million-token haystacks Kuratov et al. (2024). Although these advances surface key failure modes, they do not explore how reasoning depth interacts with very long contexts in natural prose.
@@ 3.1 Source Corpus
Hop o1 4o 4o-mini LLaMa-3.3-70B-Instruct Gemini 2.5 P Gemini 2.0 F Gemini 2.0 FL Avg. 1 95.90 95.60 92.30 94.80 96.80 93.10 90.90 94.20 2 95.50 95.40 91.80 94.40 96.50 92.80 90.30 93.81 3 95.20 95.10 91.30 94.00 96.30 92.40 90.00 93.47 4 94.80 94.90 90.90 93.60 96.20 92.10 89.60 93.16 Avg. 95.35 95.25 91.58 94.20 96.45 92.60 90.20 93.66
Table 1: Accuracy (%) of each model on NovelHopQA when evaluated with the original golden context.
Figure 1: Keyword-guided paragraph-chaining pipeline used to build NovelHopQA. See Appendix G for a full example showing multi-hop evolution across four refinement stages.
@@ 3.2 Salient Keyword Filtering
For each of the 83 novels, we prompt GPT-4o-mini OpenAI (2024a) to suggest five “anchor” keywords—characters, locations, or objects central to the plot (see Appendix I for prompt). If any keyword appears fewer than 50 times in the text, we discard and re-sample that anchor, repeating up to seven times to ensure five high-frequency anchors.
Context Hop o1 4o 4o-mini Gemini 2.5 P Gemini 2.0 F Gemini 2.0 FL LLaMa-3.3-70B-Instruct Avg. 64k 1 92.51 90.12 75.49 92.34 87.37 82.53 84.12 86.35 2 87.66↓4.85 84.25↓5.87 74.77↓0.72 87.84↓4.50 77.02↓10.35 71.39↓11.14 73.88↓10.24 79.54↓6.81 3 84.99↓2.67 81.34↓2.91 73.14↓1.63 85.12↓2.72 74.25↓2.77 70.05↓1.34 71.02↓2.86 77.13↓2.41 4 82.15↓2.84 78.47↓2.87 68.04↓5.10 82.45↓2.67 71.76↓2.49 65.33↓4.72 68.11↓2.91 73.76↓3.37 96k 1 90.35 88.83 72.25 90.12 82.26 78.44 82.04 83.47 2 85.88↓4.47 82.67↓6.16 67.44↓4.81 86.03↓4.09 74.02↓8.24 67.04↓11.40 72.33↓9.71 76.49↓6.98 3 83.41↓2.47 80.41↓2.42 66.97↓0.47 83.71↓2.32 73.38↓0.64 66.05↓0.99 68.77↓3.56 74.67↓1.82 4 80.68↓2.73 76.92↓3.91 65.59↓1.38 80.98↓2.73 70.26↓3.12 62.81↓3.24 65.95↓2.82 71.88↓2.79 128k 1 88.76 86.95 70.03 89.10 81.77 75.31 80.21 81.73 2 84.33↓4.43 80.52↓6.43 63.95↓6.08 84.70↓4.40 69.13↓12.64 62.21↓13.10 69.87↓10.34 73.53↓8.20 3 81.92↓2.41 78.03↓2.92 62.95↓1.00 82.20↓2.50 68.78↓1.35 62.07↓0.14 67.92↓1.95 71.98↓1.55 4 78.80↓3.12 74.64↓3.31 61.18↓1.77 78.55↓3.65 67.32↓1.46 57.39↓4.68 64.42↓3.50 68.90↓3.08
@@ 3.4 Multi-Hop Context Chaining & Incremental QA Generation
For each book and hop depth $H\in\{1,2,3,4\}$, we assemble contexts and QA pairs as follows (see Appendix I for all prompts):
@@
    Paragraph exclusivity: Remove selected paragraphs from the pool to prevent reuse. If no matching paragraph is found after seven attempts, abort the chain and restart with a fresh anchor.
@@ 3.5 Golden-Context Filtering
To verify answerability, we evaluate all seven models on the original golden contexts used to generate each QA pair. As shown in Table 1, all models score above 90% on average, confirming the validity of most questions. We discard any question missed by any model in the final dataset used in Section 5. Removal counts are reported in Appendix B.
@@ 3.6 Irrelevant and No-Context Sanity Check
To validate that the questions require contextual reasoning, we evaluated 800 QA pairs—100 per hop—under irrelevant and no context settings. Removing context yields low accuracies, suggesting that the tested models are typically unable to answer correctly without contextual grounding. This ensures the dataset reflects reasoning, not recall. Full results are in Appendix E.
@@ 4 Human Evaluation
Ten undergraduate validators each annotated 260 examples—40 from the 1- and 2-hop sets, and 90 from the 3- and 4-hop sets. They rated Alignment, measuring how well each QA pair matched its source context, and judged Hop Match, assessing whether the answer required exactly $H$ reasoning steps. See Appendix C for detailed results and Appendix H for the evaluation form.
@@ 5 Results and Discussion
We evaluate seven models on NovelHopQA using chain-of-thought prompts: o1 OpenAI (2024c), Gemini 2.5 Pro DeepMind (2025b), GPT-4o OpenAI (2024a), GPT-4o-mini OpenAI (2024a), LLaMA-3.3-70B-Instruct Meta (2024), Gemini 2.0 Flash, and Gemini 2.0 Flash Lite DeepMind (2025a). Table 3 summarizes model accuracy across three context lengths (64k, 96k, 128k) and four hop depths (1–4).
Impact of hop depth. All models show consistent performance drops as hop depth increases. On average, accuracy falls about 12 points from 1-hop to 4-hop at 64k. Even reasoning-focused models like Gemini 2.5 Pro and o1 decline steadily, with others typically dropping 14–17 points across hops.
Impact of context length. Longer context windows also reduce performance, though less sharply than hop depth. Most models lose about 4–6 points from 64k to 128k on 1-hop questions. This trend is especially visible among upper-mid-tier models like GPT-4o and LLaMA, which perform well under shorter contexts but degrade more under scale.
Model comparisons. Gemini 2.5 Pro and o1 consistently top each row. GPT-4o and LLaMA follow closely, both showing strong multi-hop reasoning and better robustness than smaller models. GPT-4o-mini and Flash Lite drop into the 60s under 4-hop and 128k, while Flash holds the middle but is outperformed by LLaMA at higher hops.
Robustness at scale. Despite large context windows, no model maintains strong performance on the hardest tasks (4-hop at 128k), where even top models dip below 80%. These results affirm that long-context capacity alone is not enough—multi-hop reasoning remains an open challenge. Analysis of failure modes is provided in Appendix D, with additional RAG evaluations and breakdowns included in Appendix F.
@@ 6 Conclusion
NovelHopQA is the first benchmark to vary both context length (64k–128k) and hop depth $H\in\{1,2,3,4\}$ in long-context QA. Human validation confirms quality, and models show accuracy drops along both axes. These results highlight that larger context windows aren’t enough—multi-hop reasoning remains a core challenge. We also conduct RAG evaluations F. Code, data, and more details are available at: https://novelhopqa.github.io.
@@ 8 Ethics Statement
We describe our dataset construction process in Section 3, and include all prompt templates in Appendix I. All model generations were obtained using publicly available APIs. Specifically, we used the Azure AI Foundry API for GPT-4o, GPT-4o-mini OpenAI (2024a), o1 OpenAI (2024c), and LLaMa-3.3-70B-Instruct Meta (2024); and the Google Vertex API for Gemini 2.0 Flash, Flash Lite DeepMind (2025a), and Gemini 2.5 Pro DeepMind (2025b). All models were queried using CoT prompts, and their outputs were graded with GPT-4.1 OpenAI (2024b) using CoT-based evaluation prompts. We plan to release the dataset, prompts, and model outputs upon publication to support replication and further research.
Acknowledgements
We thank Isha Sudhir, Ethan Wing, and Ernest Zhang for their helpful feedback, discussions, and support throughout the development of this work.
@@ Appendix E Irrelevant and No Context Evaluation
Interpretation.
This experiment validates the integrity of NovelHopQA by confirming that models are not simply memorizing QA pairs seen during pretraining. Accuracy remains near-zero when relevant context is removed, demonstrating that our questions are novel and context-dependent. These findings strengthen confidence that model performance on NovelHopQA reflects actual reading comprehension and not artifact exploitation or memorization.
Appendix F RAG Evaluations
Pipeline.
We divide each novel into non-overlapping chunks of 350 tokens, which are embedded using the bge-large-en encoder BAAI (2023). During inference, we use the Facebook FAISS retrieval model Research (2017) to select the top $k{=}7$ most relevant chunks based on inner-product similarity.
At inference time, we follow this process:
  1. 1.
    Encode the input question using bge-large-en.
  2. 2.
    Use the Facebook FAISS retrieval model to find the top $k{=}7$ most relevant chunks via inner-product similarity search.
  3. 3.
    Concatenate the retrieved chunks in their original order and prepend the question to form a context of roughly 2.5k tokens.
  4. 4.
    Pass this context to each of the seven models.
Why a RAG setting?
The retrieved context mimics a real-world system in which only relevant snippets—not the full 64k–128k window—are available to the generator. We expect lower accuracy because (i) retrieval can miss one or more hop paragraphs and (ii) evidence may be partial or out of order.
Results.
Table 13 reports accuracy on the same questions. As anticipated, scores cluster around 50 %, roughly 25–35 points below the golden-context setting.
Hop o1 4o 4o-mini Gemini 2.5 P Gemini 2.0 F Gemini 2.0 FL LLaMa-3.3-70B-Instruct Avg.
1 62.43 60.87 49.18 63.14 55.03 43.92 48.23 54.69
2 56.03 54.78 46.27 57.36 49.86 39.29 44.69 49.75
3 52.13 50.37 43.78 54.04 46.47 40.68 42.93 47.20
4 48.35 47.05 40.97 50.31 43.63 36.57 39.83 43.82
Avg. 54.74 53.27 45.05 56.21 48.75 40.11 43.92 48.86
Table 13: Accuracy (%) of RAG-augmented models. For each query, the retriever encoder processed the full novel to generate vector embeddings for all chunks, which were then used to retrieve the top $k{=}7$ most relevant chunks.
Analysis.
Despite feeding models the full retrieved context directly in the input prompt, LLMs still underperform.. This gap underscores the limitations of retrieval-augmented reasoning even when relevant evidence is available at inference time. Our error analysis identifies two dominant failure modes:
  1. 1.
    Incomplete hop coverage — For multi-hop questions, crucial context is often fragmented across the book. When even one necessary chunk is missing from the top-$k$ retrieved segments, models fail to complete the reasoning chain. This issue becomes more pronounced in 3- and 4-hop examples, where partial evidence leads to confidently incorrect answers.
  2. 2.
    Evidence misalignment — Unlike the golden oracle where evidence is presented in coherent order, RAG-retrieved chunks may be out of narrative sequence or lack discourse continuity. This can confuse even strong models, leading to misinterpretation of character arcs, event timelines, or causal links.
These failure modes point to the brittleness of current RAG pipelines when applied to deep reasoning tasks over long texts. They support our broader conclusion: retrieval helps, but cannot replace full-context comprehension in complex multi-hop reasoning. Addressing this limitation may require more sophisticated retrieval strategies—such as iterative chunk expansion, retrieval-conditioned generation, or train-time exposure to fragmented narratives.
F.1 Visual Breakdown of RAG Accuracy
Figure 5: Left: Heatmap of RAG-augmented accuracy across hop depths for all models. Right: Accuracy curves for each model as a function of hop level. Together, these plots show how performance degrades with deeper reasoning steps even when retrieval narrows the context.
Appendix G 4-Hop QA Evolution Example
@@ F.1 Visual Breakdown of RAG Accuracy
Appendix H Human Evaluation Form Example
@@
Appendix I Prompt Templates